# Representational Alignment Supports Effective Machine Teaching

### v1

@@ B.3 Utility Curve over Classrooms
The utility curves that we construct in Section 3 and B.2 were always constructed with respect to a single student (in a respective, “dyadic” [^3] Pairing two agents – one student and one teacher.). However, in our classroom settings, we also corrupt the students’ representations to simulate representational diversity. We sample a new utility curve, wherein, for each teacher parameterization (same error rate parameterization as above, with representation corruptions now in increments of $0.1$), we sample $10$ different student corruptions ranging over 00 to $0.9$ in increments of $0.1$. We build this curve only for the column label type. We then bucketize the teacher error rate as well as the representation alignment such that we can index into the curve to extract an “expected average performance” for any student-teacher pair.
B.4 Constructing Classroom Pools
@@
In the structured setting, we construct clusters of similar students and teachers. We prespecify a number of clusters $M$ and number of students per cluster. Clusters are designed to span a range of levels of representation alignment over the “original” grid. We loop over possible representation alignments corruptions ranging from 00 to $1$ in increments of $1/M$. For each cluster, we sample a “seed” student using that corruption level. We then sample [notation?] students ontop of this cluster with a representation corruption of $0.01$ ontop of the base student to ensure students share similar (but some variation) in their representation. For each cluster, we sample a teacher with error rate uniformly from $0-0.5$ and representation with a similar slight possible corruption (sampled uniformly from $0-0.01$) ontop of the seed student, thereby ensuring that there would be a representationally similar teacher for each student in each cluster if provided. However, to simulate gaps in coverage of particular representation characterezations, we randomly drop some teachers from the pool.
B.5 Generalization to the Dino Stimuli
@@
Table 4: Student learning outcomes (accuracy) from different classroom matching approaches in the unstructured pool setting for the dino stimuli. We again compute the student performance across all $N=1000$ pooled students (paired with potentially 30 teachers). Error bars are again computed over 10 different sampled pools.
@@
Random 0.29 $\pm$ 0.00 0.16 $\pm$ 0.01 0.46 $\pm$ 0.01 0.04 $\pm$ 0.01
MOOC 0.34 $\pm$ 0.02 0.22 $\pm$ 0.00 0.51 $\pm$ 0.07 0.09 $\pm$ 0.05
Ours 0.36 $\pm$ 0.01 0.22 $\pm$ 0.00 0.64 $\pm$ 0.08 0.15 $\pm$ 0.03
Optimal 0.44 $\pm$ 0.01 0.31 $\pm$ 0.00 0.69 $\pm$ 0.08 0.32 $\pm$ 0.02
@@
Random 0.29 $\pm$ 0.02 0.18 $\pm$ 0.02 0.44 $\pm$ 0.03 0.06 $\pm$ 0.03
MOOC 0.34 $\pm$ 0.03 0.23 $\pm$ 0.02 0.50 $\pm$ 0.08 0.10 $\pm$ 0.06
Ours 0.35 $\pm$ 0.03 0.23 $\pm$ 0.02 0.54 $\pm$ 0.08 0.13 $\pm$ 0.06
Optimal 0.38 $\pm$ 0.02 0.28 $\pm$ 0.01 0.56 $\pm$ 0.07 0.15 $\pm$ 0.06
B.6 Additional Details on Student-Centric Teacher
In contrast to our self-centered teacher, our student-centric teacher does not use its own representation to select examples to provide to the student. Instead, the student-centric teacher is endowed with an inner optimization loop over the students assigned to it, whereby the teacher loops $T$ times over the students and randomly selects one point per category (using the teacher’s believed class – the teacher may not know the true categories) and measures the expected performance of each student if that set of examples were revealed. Note, the teacher computes the expected accuracy of each student using against its belief of the true categorization (which may be incorrect). The teacher then chooses the set of examples that attains the highest average accuracy over students. Here, we set $T$ to 100; exploring the impact of varied $T$ is a sensible next step. Exploration of alternate optimization functions, e.g., optimizing over the minimum attained performance over the students in the teacher’s classroom rather than average classroom performance are also ripe ground for future work.
B.7 Relationship Between Classroom Size and Student-Centric Teacher Utility
We conduct a deeper dive into the relationship between classroom size and student performance by sampling sweeping over a range of teacher error rates (between 00 and $0.5$ in increments of $0.1$) and classroom sizes. We sweep over 10 samples for each setting, where each self-centered teacher optimizes for their pool through $T=100$ inner loop iterations. Students are drawn from our structured, clustered pool. We then marginalize over our sampled error rates to compute an expected average classroom accuracy per classroom size. We find that the performance of students in a classroom with a student-centric teacher falls off rapidly as a function of classroom size and then plateaus in Figure 7.
Figure 7: Average performance of students paired with a student-centric teacher as a function of classroom size. Error bars depict standard error over runs.
Appendix C Additional Human Experiment Details
C.1 Participant Recruitment and Compensation

### v2

@@ F.2 Utility curve over classrooms
The utility curves that we construct in Section 2 and E.2 were always constructed with respect to a single student (in a respective, “dyadic” [^3] Pairing two agents – one student and one teacher.). However, in our classroom settings, we also corrupt the students’ representations to simulate representational diversity. We sample a new utility curve, wherein, for each teacher parameterization (same error rate parameterization as above, with representation corruptions now in increments of $0.1$), we sample $10$ different student corruptions ranging over 00 to $0.9$ in increments of $0.1$. We build this curve only for the column label type. We then bucketize the teacher error rate as well as the representation alignment such that we can index into the curve to extract an “expected average performance” for any student-teacher pair.
F.3 Constructing classroom pools
@@
In the structured setting, we construct clusters of similar students and teachers. We prespecify a number of clusters $M$ and number of students per cluster. Clusters are designed to span a range of levels of representation alignment over the “original” grid. We loop over possible representation alignments corruptions ranging from 00 to $1$ in increments of $1/M$. For each cluster, we sample a “seed” student using that corruption level. We then sample students on top of this cluster with a representation corruption of $0.01$ on top of the base student to ensure students share similar (but some variation) in their representation. For each cluster, we sample a teacher with error rate uniformly from $0-0.5$ and representation with a similar slight possible corruption (sampled uniformly from $0-0.01$) on top of the seed student, thereby ensuring that there would be a representationally similar teacher for each student in each cluster if provided. However, to simulate gaps in coverage of particular representation characterizations, we randomly drop some teachers from the pool.
F.4 Additional classroom matching results
We include additional results into classroom matching in Tables 2 and 3 and a relationship between group size and learning outcomes in Figure  5.
Table 2: Student learning outcomes (accuracy) from different classroom matching approaches in the structured pool setting. Higher is better for all metrics. $\pm$ indicates standard errors computed over 40 sampled pools and associated assignments. We compute the average student performance across all $N=1000$ pooled students (paired with potentially $M=30$ teachers), as well as accuracy over the bottom and top 10% of students in each matching, respectively. We additionally compute the proportion of students who achieve “passing” marks (set to a moderately high threshold of 45% accurate, given chance guessing is 16.6% on our 6x6 grid). Higher is better for all metrics. $\pm$ indicates standard errors computed over 10 sampled pools and associated assignments.
Method Avg Acc Bottom 10% Top 10% Pass Rate
Random 0.33 $\pm$ 0.01 0.21 $\pm$ 0.01 0.49 $\pm$ 0.01 0.12 $\pm$ 0.02
Min Err 0.37 $\pm$ 0.01 0.26 $\pm$ 0.01 0.53 $\pm$ 0.03 0.18 $\pm$ 0.04
Utility 0.38 $\pm$ 0.01 0.27 $\pm$ 0.01 0.55 $\pm$ 0.03 0.20 $\pm$ 0.04
Optimal 0.43 $\pm$ 0.01 0.32 $\pm$ 0.01 0.60 $\pm$ 0.02 0.32 $\pm$ 0.04
Table 3: Student learning outcomes (accuracy) from different classroom matching approaches in the unstructured pool setting. $\pm$ indicates standard errors computed over 40 sampled pools and associated assignments.
Method Avg Acc Bottom 10% Top 10% Pass Rate
Random 0.33 $\pm$ 0.00 0.17 $\pm$ 0.01 0.52 $\pm$ 0.01 0.09 $\pm$ 0.01
Min Err 0.39 $\pm$ 0.01 0.25 $\pm$ 0.00 0.57 $\pm$ 0.02 0.20 $\pm$ 0.02
Utility 0.39 $\pm$ 0.01 0.24 $\pm$ 0.00 0.61 $\pm$ 0.02 0.23 $\pm$ 0.02
Optimal 0.49 $\pm$ 0.00 0.36 $\pm$ 0.00 0.71 $\pm$ 0.02 0.54 $\pm$ 0.01
F.5 Generalization to the dino stimuli
@@
Table 4: Student learning outcomes (accuracy) from different classroom matching approaches in the unstructured pool setting for the dino stimuli. We again compute the student performance across all $N=1000$ pooled students (paired with potentially 30 teachers). Error bars are again computed over 40 different sampled pools.
@@
Random 0.29 $\pm$ 0.00 0.16 $\pm$ 0.00 0.47 $\pm$ 0.01 0.04 $\pm$ 0.00
Min Err 0.35 $\pm$ 0.01 0.22 $\pm$ 0.00 0.55 $\pm$ 0.04 0.13 $\pm$ 0.03
Utility 0.36 $\pm$ 0.01 0.22 $\pm$ 0.00 0.62 $\pm$ 0.04 0.16 $\pm$ 0.02
Optimal 0.44 $\pm$ 0.01 0.31 $\pm$ 0.00 0.69 $\pm$ 0.03 0.33 $\pm$ 0.01
@@
Random 0.30 $\pm$ 0.01 0.18 $\pm$ 0.01 0.45 $\pm$ 0.02 0.06 $\pm$ 0.01
Min Err 0.34 $\pm$ 0.01 0.23 $\pm$ 0.01 0.51 $\pm$ 0.05 0.10 $\pm$ 0.03
Utility 0.36 $\pm$ 0.01 0.23 $\pm$ 0.01 0.57 $\pm$ 0.05 0.15 $\pm$ 0.03
Optimal 0.39 $\pm$ 0.01 0.28 $\pm$ 0.01 0.59 $\pm$ 0.04 0.18 $\pm$ 0.03
F.6 Additional details on student-centric teacher
In contrast to our self-centered teacher, our student-centric teacher does not use its own representation to select examples to provide to the student. Instead, the student-centric teacher is endowed with an inner optimization loop over the students assigned to it, whereby the teacher loops $T$ times over “simulated students” (which we call the “inner loop”) and randomly selects one point per category (using the teacher’s believed class – the teacher may not know the true categories) and measures the expected performance of each student if that set of examples were revealed. Note, the teacher computes the expected accuracy of each student using against its belief of the true categorization (which may be incorrect). The teacher then chooses the set of examples that attains the highest average accuracy over students. Here, we set $T$ to 100; exploring the impact of varied $T$ is a sensible next step. Exploration of alternate optimization functions, e.g., optimizing over the minimum attained performance over the students in the teacher’s classroom rather than average classroom performance, as well as exploring different kinds of simulated students (here, we assume the teacher’s have the right model of each student) are also ripe ground for future work.
We explore the effect of student-centric teachers by appending a second stage to our matching procedure. After matching using our utility curve (as noted above), we greedily attempt to pair the lowest performing students with a student-centric teacher who chooses points by optimizing for the students in their pool (i.e., taking the students’ representations into account). We continue incorporating the next lowest-performing students into the student-centric teacher’s classroom until a student’s attained accuracy with the original pairing is not improved by the student-centric teacher. We apply our procedure to the clustered pool structure noted above and find that it is beneficial to continue adding students up to a point: if the teacher is an expert (zero error rate), we can add all students from one cluster before we see detrimental performance across the pool of students assigned to said teacher. As the student-centric teacher’s error rate increases, fewer students can be pooled before performance dropoff (see Appendix Figure 5).
These results indicate the student-centric teachers can cover students who are representationally distinct and help boost their learning outcomes. However, classroom size matters, corroborating prior works in machine and human teaching (Frank, 2014; Yeo et al., 2019; Ma et al., 2018; Zhu et al., 2018). In the next section of the Appendix, we conduct a deeper dive into the relationship between classroom size and student outcomes in our setting when student-centric teachers are available. Herein, we see that teachers who may try to overalign to all students at once in a large classroom induce poorer outcomes for the classroom writ large.
Appendix G Additional human experiment details
G.1 Participant recruitment and compensation